OpenStudy (anonymous):

Find the percent of each number. 3.5% of 700

OpenStudy (aroub):

just simplify it! I'm sure you know how to do it on your own right?

OpenStudy (anonymous):

3.5 * 700 --- 100

jhonyy9 (jhonyy9):

3,5 ---- 100 x ----- 700 ------------- x=(3,5*700)/100 = 3,5*7=24,5